b. A rectangular water tank has dimensions of 2.0 m x 1.0 m x 1.0 m. What would be the diameter of a 1.50-m high cylinder tank to contain the same amount of water that the rectangular tank contains
1
Expert's answer
2020-10-05T10:54:08-0400

The volume of a rectangular water tank

"V=a\times b\times c"

The volume of a cylinder water tank

"V=\pi R^2h"

Hence, the high of cylinder tank

"h=\frac{abc}{\pi R^2}=\rm \frac{2.0\times 1.0\times 1.0}{3.14\times (1.50/2)^2}=1.13\: m"